OpenStudy (anonymous):

does anyone understand how to solve quadric equations how factoring is used to solve quadratic equations. Demonstrate the process with an example.

OpenStudy (anonymous):

The general form of the quadratic eqn is ax^2 + bx +c

OpenStudy (anonymous):

We first take the product of a and c , keeping in mind their + or - sign

OpenStudy (anonymous):

then we find two numbers, say p and q, such that p x q = ac and p + q = b

OpenStudy (anonymous):

Then we substitute bx by px + qx in the eqn and then find factors

OpenStudy (anonymous):

Understood till now???

OpenStudy (anonymous):

for eg 2x^2 - 4x - 6 Here a=2 , b= -4 and c = -6

OpenStudy (anonymous):

Here a x c = 2 x -6 = -12 Now we have to break up -12 so that the two factors give product -12 and sum -4 Factors of 12 can be 1 x 12, 2 x 6, 3 x 4, Out of these, if we take the pair 2 x 6 and then modify it as 2 x -6, we get 2 x -6 = -12 and 2 + (-6) = -4 So, we can say 2x^2 -4x -6 = 2x^2 + 2x - 6x - 6 = 2x(x + 1) -6 (x + 1) Hence 2x^2 -4x -6 = (2x - 6) (x + 1)

OpenStudy (anonymous):

Hope this clears up factorisation for you!!